The following table presents the observed and expected data on the number of plants found in each of 50 sampling quadrants.

# of plants Observed Frequency (Oi) Expected Frequency (Ei)

0

5 6.676
1 18 13.534
2 10 13.534
3 12 9.022
\geq4 5 7.144

H0 : The distribution is Poisson

H1 : The distribution is not Poisson

a.) Justify why the assumption of the Poisson distribution seem appropriate as a probability model for this data? Find the value of the only parameter Poisson distribution has.

b.) Using α = 0.01, find the test statistic and critical region to make a conclusion about this goodness of fit test. (Hint: we did not estimate the parameter)

c.) Calculate the P-value for this test. State the conclusion for this test.
Answer #1

Solve using Excel:

D E F 1 Expected Probability (0-5) / E 2 3 А B C No. of Plants (x) | Observed Frequency (0) x*Observed Frequency 0 5 0.00 1 1

No. of Plants (x) Observed Frequency (O) x*Observed Frequency Expected Probability Expected Frequency (E) (Oi - Ei)2 / Ei
0 5 0.00 0.15 7.63 0.91
1 18 18.00 0.29 14.34 0.93
2 10 20.00 0.27 13.48 0.90
3 12 36.00 0.17 8.45 1.49
4 or more 5 20.00 0.12 6.09 0.20
Total 50 94.00 Total 4.43
λ 1.88
No. of Plants (x) Observed Frequency (O) x*Observed Frequency Probability = nCx*px*qn-x Expected Frequency (E) (Oi - Ei)2 / Ei
0 50 =A2*B2 =POISSON.DIST(A2,$B$8,FALSE) =D2*50 =((B2-E2)^2)/E2
1 19 =A3*B3 =POISSON.DIST(A3,$B$8,FALSE) =D3*50 =((B3-E3)^2)/E3
2 7 =A4*B4 =POISSON.DIST(A4,$B$8,FALSE) =D4*50 =((B4-E4)^2)/E4
3 4 =A5*B5 =POISSON.DIST(A5,$B$8,FALSE) =D5*50 =((B5-E5)^2)/E5
4 or more 0 =4*B6 =1-SUM(D2:D5) =D6*50 =((B6-E6)^2)/E6
Total =SUM(B2:B6) =SUM(C2:C6) Total =SUM(F2:F6)
λ =C7/B7

H0 : The distribution is Poisson

H1 : The distribution is not Poisson

a) Only parameter: λ = 1.88

b) Test statistic = 4.43

Level of significance = 0.01

Degrees of freedom: df = n-k-1 (k is no. of parameters) = 50-1-1 = 48

Critical value (Using Excel function CHISQ.INV.RT(probability,df)) = CHISQ.INV.RT(0.01,48) = 73.68

c) p-value (Using Excel function CHISQ.DIST.RT(test statistic,df)) = CHISQ.DIST.RT(4.43,48) = 1

Since p-value is more than 0.01, we do not reject the null hypothesis.

So, the distribution is Poisson.